# Scalable Quantum Simulations of Scattering in Scalar Field Theory on 120 Qubits

### v1

@@ IV Quantum simulation of scalar field theory using scalable variational circuits
The scalable variational circuits procedure can be carried out on a classical computer or a small quantum computer. Throughout this work, classical computational resources are leveraged to determine parameters that are extrapolated to system sizes of interest to be run on quantum computers. This work uses $n_{q}=2$ qubits to represent the state of $\phi$ at each site, for the theory with bare mass $m=1/2$. The method of Ref. [82] is used to choose the cutoff $\phi_{\text{max}}=1.5$ to minimize field digitization effects. The dynamics of wavepackets in the free $(\lambda=0)$ theory is compared to those in the interacting theory with coupling $\lambda=2$. For the parameters chosen, the correlation lengths $\xi_{\lambda}$ of the free and interacting theory are $\xi_{0}=2.4503$ and $\xi_{2}=1.5817$, respectively. These are determined, as in the previous section, by computing the inverse gap for small but increasing system sizes $L$, fitting an exponential, and extrapolating to the infinite-volume limit. The circuit elements that are used to compose the variational circuits throughout this section are shown in Fig. 3. Details of the optimization strategies and choices are given in App. B. The resulting parameters $\vec{\theta}_{\text{opt}}$ for all components of the variational circuit are given in App. D.
@@ IV.1 Vacuum preparation
In the free theory, $|\psi_{\text{vac}}\rangle$ defined in Eq. (7) can be used as $|\psi_{\text{targ}}\rangle$. Approximations to the vacuum of the interacting theory can be found analytically using standard perturbative methods for small $\lambda$ as is discussed in Sec. II. Fortunately, for the chosen system parameters $m$ and $\lambda$, the variational parameters of the circuits initializing these states show convergence for small-enough system sizes where exact diagonalization (ED) is viable. Taking advantage of this, interacting ground state determined from ED is used as $|\psi_{\text{targ}}\rangle$.
@@
The Hamiltonian of the theory is real, so it must have real eigenstates. By using exponentials of commutators of terms in the Hamiltonian for the operator pool, SC-ADAPT-VQE ensures that only real unitaries are used, and maintains the symmetries of the system. The lowest order operator in $\phi$ and $\Pi$ that acts over more than one site and is not included in the Hamiltonian is derived from the commutator of terms in $H_{\text{kin}}$ and $H_{\Pi}$, $O_{1}=\frac{1}{2}\sum_{j}[\phi_{j}\phi_{j+1},\Pi^{2}_{j}]$. Because of the chosen form of $\Pi_{j}$, the canonical commutation relation $[\phi_{i},\Pi_{j}]=i\delta_{ij}$ only holds in the limit $\delta_{\phi}\rightarrow 0$ (i.e., only for the undigitized operators $\phi_{j}$, $\Pi_{j}$). Nevertheless, this relation can still be used at nonzero $\delta_{\phi}$ to simplify the operators in the SC-ADAPT-VQE pool because the requisite symmetries are maintained. This makes operators such as $O_{1}$ easier to implement using quantum gates. Using this, $O_{1}$ can be rewritten as
$O_{1}=i\sum_{i}\Pi_{i}\phi_{i+1},$ (22)
@@
$O_{d}=i\sum_{i}\Pi_{i}\phi_{i+d}.$ (23)
This operator can similarly be obtained from the commutator of $\Pi_{j}^{2}$ with a longer-range verion of $H_{\text{kin}}$. The pool used for SC-ADAPT-VQE consists of the operators $\{O_{d}\}$ for a range of increasing $d$.
@@
Figure 4 shows the convergence of the variationally-prepared vacuum. The local infidelity over four sites, $I_{4}$, is used to measure the quality of prepared state, spanning roughly 2-3$\times$ the correlation length of the infinite-volume vacuum. With just one SC-ADAPT-VQE layer, $I_{4}$ can be seen to be non-increasing in Fig. 4a, meaning that the local quality of the prepared function is not degrading with system size for both $\lambda=0$ and $\lambda=2$. The infidelity is systematically improved by applying more layers of $O_{d}$ with successively larger $d$. The result of longer-range operators creating correlations that span more lattice sites is seen in the plots of $I_{4}$ in Fig. 4b. In both of these figures, it can be seen that the $\lambda=2$ wavefunction is converging faster as a result of the smaller correlation length in the interacting theory. Figure 4c is a representative example of the convergence of the parameters with increasing system size. These plots indicate that the $L$-extrapolation is applicable and the parameters to initialize the ground state for an arbitrary system size can be determined using modest $L$, where classical simulation is possible. [^12] If the SVC method is run using classical computation, the correlation length of ground states that can be prepared with the presented method is limited by the maximum number of qubits that can be simulated classically.
Figure 4: Convergence of the wavefunction and state preparation parameters in the vacuum preparation circuit shown in Fig. 3b for both the free and interacting theory. (a): $I_{4}$ taken over a region of four contiguous lattice sites (eight qubits) as a function of system size. One layer of SC-ADAPT-VQE is used. (b): $I_{4}$ as a function of number of SC-ADAPT-VQE layers used in the ansatz for a $L=8$ system. Each layer uses a longer-range operator $O_{d}$ given in Eq. 23 to create correlations that span more lattice sites. (c): An example of the exponential convergence of one of the angles parameterizing the $\lambda=0$ vacuum preparation circuit with one layer of SC-ADAPT-VQE, as a function of system size. The remainder of the angles are given numerically in App. D.
@@ IV.2 Wavepacket creation
The creation of particles on top of the prepared vacuum requires initializing local excited states. JLP proposed a method of creating wavepackets in the free theory by applying unitary evolution under $a_{k}^{\dagger}$ and $a_{k}$ defined in Eq. (3), and then adiabatically turning on the interaction as described in Sec. II. Instead of this, SVC is used to compress the wavepacket preparation circuit by taking advantage of classical computation. This step of the algorithm prepares a wavepacket whose size remains fixed as the total system size increases. While this method does not allow for the scalable determination of the variational parameters to prepare wavepackets of increasing size, it utilizes the scalability in total system size, including vacuum regions. [^13] In this case, the classical computing capability limits the maximum size of the wavepacket. This is useful because a finer spacing in $k$-space increases the spatial momentum resolution of the wavepacket, building a more accurate representation of wavepackets in the continuum.
@@
These are plotted as the dashed and faint solid lines respectively in Fig. 5b. $E_{k}$ and $v_{k}$ can be computed numerically in the digitized theory for any value of $\lambda$ by using ED and projecting onto each momentum sector. These numerically-calculated values are shown in Figs. 5a and b by the solid lines with circle and square markers.
@@
Figure 5: (a): The dispersion relation $E_{k}$ and (b): the group velocity $v_{k}$ for $m=1/2$ and $\lambda=0,2$. Exact computation is possible in the continuum (faint solid lines) and in the free lattice theory (dashed lines). By exactly diagonalizing the Hamiltonian of Eq. (1) in the single-particle sector at a fixed lattice size, it is possible to keep track of the lowest eigenvalues as a function of $\lambda$, matching them onto the exact values at $\lambda=0$ to construct $E_{k}$ and $v_{k}$ numerically (solid lines with circle and square markers). The numerical computations are done on a lattice of $L=8$ spatial sites with $n_{q}=3$ qubits per site. Discrete gradients are used for the numerical computations of $v_{k}$. (c): The convergence of the prepared particle wavepackets occupying three spatial sites with momentum $k=-\pi/3$, as a function of number of layers of the circuit from Fig. 3c. The local infidelity spanning the width of the wavepacket, $I_{3}$, is used to measure the quality of the prepared wavefunction for $\lambda=0,2$. Compared to (a) and (b), (c) uses $n_{q}=2$ qubits to represent the field at each site.
[…]
@@
Leveraging the analytic knowledge of the free theory, the wavefunction of the wavepacket $|\psi_{\text{wp}}\rangle$ defined in Eq. (10) is used as $|\psi_{\text{targ}}\rangle$ for the variational optimization. The adiabatic turn-on of interactions given by Eq. (12) is simulated classically, producing interacting wavepacket wavefunctions that are then used for $|\psi_{\text{targ}}\rangle$ in SVC for the interacting theory. With this method, interacting wavepackets are prepared directly on top of the interacting vacuum. This way, in addition to reducing the circuit depth, SVC also removes the necessity for implementing the adiabatic evolution to turn on interactions.
@@
Because the region where the desired wavepacket has support does not grow with increasing system size $L$, the amount of qubits over which the variational circuit has to act is fixed by the spatial width of the wavepacket. Scaling these circuits is simple and straightforward since the number of parameters and structure do not change with system size. This is another example of using physical properties of the state, in this case locality of the wavepacket, to guide the design of the variational circuit. One layer of the brickwall ansatz used for this step of the state preparation is shown in Fig. 3c. The building blocks of this circuit are the variational gates defined in Ref. [133]. This block is particularly useful because any product of gates from the universal gate set $\{\text{CNOT},R_{X},R_{Y},R_{Z}\}$ can be written in terms of products of the block with different parameters. Each layer of the brickwall ansatz has a two-qubit gate depth of 2 and uses 20 variational parameters.
@@
Figure 5c shows the convergence of the prepared wavepacket state as a function of the number of layers of the brickwall ansatz of Fig. 3c. It is seen that the local infidelity spanning the wavepacket width, $I_{3}$, is systematically improved by adding more layers. Four layers of the circuit shown in Fig. 3c are used to prepare the free and interacting wavepackets. [^15] Note that with the parameters set to zero, inserting CNOTs between layers of this circuit will collapse the circuit to the identity. This is used in vacuum simulations to maintain the structure of the circuit. These brickwall circuits have two-qubit gate depths of 14 (including extra site-wise SWAP gates), and 80 parameters in total. While it is possible to run an extrapolation of the wavepacket parameters, this is not necessary. The angles parameterizing the vacuum creation circuits are observed to change on the scale of $10^{-4}-10^{-3}$ for $L=10-12$ (see Table 5), which is at the precision tolerance of IBM devices. The extrapolated angles for the vacuum preparation are used to prepare the equivalent vacuum state on a smaller system size for purposes of training the wavepacket circuit. The vacuum preparation angles extrapolated to $L=60$ are used in an $L=12$ system to optimize the wavepacket preparation parameters. The resulting parameters $\vec{\theta}_{\text{opt}}$ are given in Table 6, and numerical $I_{3}$ values are given in Table 1.
@@ IV.3 Time evolution
In this work, a translationally-invariant brickwall ansatz is used to compress the time evolution operator. Note that while there is no guarantee that a shallow brickwall circuit can approximate the time evolution operator well, time evolution under any Hamiltonian with local interactions is subject to Lieb-Robinson bounds [150]. Moreover, there is a maximum value for the group velocity of excitations from the lattice dispersion relation (Eq. (24)) [51] so it is reasonable to expect that low-depth brickwall circuits can reliably compress evolution under the Hamiltonian of Eq. (1) for sufficiently early times. The ansatz is made up of layers of the circuit shown in Fig. 3e, arranged in a symmetric way, seen in Fig. 6. Each variational step is made up of six forward layers and six reverse layers, with the reverse layers having the same parameters as the forward layers. This structure is chosen to mimic the structure of a second-order Trotter step. Each layer uses 12 parameters and has a two-qubit gate depth of two. The time evolution of the two-wavepacket state $e^{-itH}|\psi_{\text{2wp}}\rangle$ on an $L=12$ system is used as $|\psi_{\text{targ}}\rangle$ to determine the variational circuits.
@@
Similar to the wavepacket preparation step, the $L$-extrapolation is not necessary for time evolution. However, in this case this simplification is only valid for early times, because the time evolution circuits act on the whole lattice. Since the vacuum is not prepared exactly, and the time evolution is not implemented exactly, the vacuum regions will evolve in time [20]. This evolution will be captured by the variational optimization procedure, up to finite-size effects that scale as $O(t/L)$. [^16] These effects are estimated and the vacuum evolution is used for error mitigation. See App. C for details. To get around this limitation, the optimization can be run on larger systems for later times, or an $L$-extrapolation can be implemented. Furthermore, propagating particles in a scattering simulation show more drastic effects from finite-size effects, such as wrapping around the lattice due to PBCs and re-scattering. This also necessitates running the optimization for later simulation times on a larger lattice. While this is a limitation of the method in general, the scattering simulations are set up to require minimal time before scattering is observed (see Sec. IV.2 for details). This places a limit both on the time required in the simulation, as well as the size of the interaction region that needs to be used for training the variational circuits.
@@
One promising way to use variational circuits for time evolution without running into these finite-size effects is to compress a single time evolution step, and use that step repeatedly in place of a Trotter step. Similar methods are investigated in Refs. [99, 104]. While the resulting circuits will be less compressed than those produced by the current method, they will be more scalable with $L$. Another interesting direction is using MPS simulations of the time evolution operator for training [144, 104, 147]. This would remove the limitations placed on this method by finite-size effects. In this setup, the time evolution of one or many states may be used for training, as described in Sec. III. Methods that target time evolution operators in specific subspaces, such as restricted to the low-energy subspace, are currently being developed [145, 108]. The combination of subspace methods with variational time evolution has the potential to simplify the optimization task while also improving the quality of the produced circuits.
@@
Figure 7 shows the convergence of the time evolution variational circuits, compared to second-order Trotter circuits as a function of circuit depth. The plots show the infidelity as a function of two-qubit gate depth for the time evolution of a single wavepacket on an $L=6$ system. The Trotter time evolution circuits are constructed by adding more Trotter steps as described in Sec. II.1. Variational circuits fix the number of variational steps to two while increasing the depth of each step by adding layers of the circuit elements in Fig. 3d following the structure of Fig. 6. The results for $t=1-9$ show that the brickwall circuits are able to consistently compress the Trotter circuits. This can especially be seen at late times, where considerably deeper Trotter circuits are required to accurately implement the time evolution operator, whereas the variational circuits show a modest growth in depth. Sharp dips in the curves of both the variational and Trotter infidelities are results of finite-size effects, as well as of numerical optimization (for the variational lines). A difference in the infidelity between $\lambda=0$ (top row) and $\lambda=2$ (bottom row) can be seen, which is attributed to using fixed-depth circuits to describe states with different correlation lengths. These plots show that by increasing the number of layers (and so, the number of parameters), the variational ansatz is able to represent the time-evolved state with increasing accuracy.
@@
Figure 7: The convergence of the time-evolved single-wavepacket state on an $L=6$ system with $\lambda=0$ (top row) and $\lambda=2$ (bottom row) for $t=1-9$. The second-order Trotter time evolution (circles) is compared to the variational steps (squares) as a function of two-qubit gate depth. Two variational steps consisting of an increasing number of layers shown in Fig. 3d are used. The infidelity of the whole $L=6$ system with the exact time-evolved state is used to measure the quality of the wavepacket state as it propagates through the lattice.
@@
For the runs on the device, the number of variational steps is increased with the simulation time: $t=\{1,2,3\}$ use one variational step, $t=\{4,5,6\}$ use two, and $t=\{7,8,9\}$ use three. The number of layers in each step is kept fixed to six forward and six reverse layers, as in Fig. 6. The steps are optimized separately for each simulation time to find better quality solutions for the variational circuit. They are applied in a Trotter-like fashion, so that the number of parameters is the same for all time steps. The time evolution portion of the variational circuit uses 72 parameters in total. The parameters $\vec{\theta}_{\text{opt}}$ that implement the time evolution are given in Tables 7-15 and the local infidelities $I_{10}$ for each time step are shown in Table 1.

### v2

@@ IV Quantum simulation of scalar field theory using scalable variational circuits
The scalable variational circuits procedure can be carried out on a classical computer or a small quantum computer. Throughout this work, classical computational resources are leveraged to determine parameters that are extrapolated to system sizes of interest to be run on quantum computers. This work uses $n_{q}=2$ qubits to represent the state of $\phi$ at each site, for the theory with bare mass $m=1/2$. The method of Ref. [85] is used to choose the cutoff $\phi_{\text{max}}=1.5$ to minimize field digitization effects, which are studied in App. A. The dynamics of free $(\lambda=0)$ and interacting $(\lambda=2)$ wavepackets is compared. For the parameters chosen, the correlation lengths $\xi_{\lambda}$ of the free and interacting theory are $\xi_{0}=2.4503$ and $\xi_{2}=1.5817$, respectively. These are determined, as in the previous section, by computing the inverse gap for small but increasing system sizes $L$, fitting an exponential, and extrapolating to the infinite-volume limit. The circuit elements that are used to compose the variational circuits throughout this section are shown in Fig. 3. Details of the optimization strategies and choices are given in App. D. The resulting parameters $\vec{\theta}_{\text{opt}}$ for all components of the variational circuit are given in App. F.
@@ IV.1 Vacuum preparation
In the free theory, $|\psi_{\text{vac}}\rangle$ defined in Eq. (7) can be used as $|\psi_{\text{targ}}\rangle$. Approximations to the vacuum of the interacting theory can be found analytically using standard perturbative methods for small $\lambda$ as is discussed in Sec. II. Fortunately, for the chosen system parameters $m$ and $\lambda$, the variational parameters of the circuits initializing these states show convergence for small-enough system sizes where exact diagonalization (ED) is viable. Taking advantage of this, the interacting ground state determined from ED is used as $|\psi_{\text{targ}}\rangle$.
@@
The Hamiltonian of the theory is real and time reversal invariant, so it must have real eigenstates. It is found that an operator pool built out of commutators of terms in the Hamiltonian, similar to the pools used in Refs. [10, 15], is effective. A pool with this structure ensures the reality of the prepared vacuum by applying only real unitaries, and maintains the symmetries of the system. The lowest-order operator in $\phi$ and $\Pi$ that acts over more than one site and is not included in the Hamiltonian is derived from the commutator of terms in $H_{\text{kin}}$ and $H_{\Pi}$, $O_{1}=-\frac{i}{2}\sum_{j}[\phi_{j}\phi_{j+1},\Pi^{2}_{j}]$. The canonical commutation relation $\phi_{i},\Pi_{j}]=i\delta_{ij}$ only holds in the limit $\delta_{\phi}\rightarrow 0$ (i.e., only for the undigitized operators $\phi_{j}$, $\Pi_{j}$). Nevertheless, this relation can still be used at nonzero $\delta_{\phi}$ to simplify the operators in the SC-ADAPT-VQE pool because the requisite symmetries are maintained. This makes operators such as $O_{1}$ easier to implement using quantum gates. Using this, $O_{1}$ can be rewritten as
$O_{1}=\sum_{j}\Pi_{j}\phi_{j+1},$ (22)
@@
$O_{d}=\sum_{j}\Pi_{j}\phi_{j+d}.$ (23)
This operator can similarly be obtained from the commutator of $\Pi_{j}^{2}$ with a longer-range verion of $H_{\text{kin}}$. The pool used for SC-ADAPT-VQE consists of the operators $\{O_{d}\}$ for a range of increasing $d$. Note that the operators $O_{d}$ are imaginary in the $\phi$-basis. This is because $\phi_{j}$ is real and $\Pi_{j}$ is imaginary since it is the conjugate momentum operator. As a result, $e^{-i\theta O_{d}}$ is a real unitary.
@@
[…]
Figure 4: Convergence of the wavefunction and state preparation parameters in the vacuum preparation circuit shown in Fig. 3b for both the free and interacting theory. (a): $I_{4}$ taken over a region of four contiguous lattice sites (eight qubits) as a function of system size. One layer of SC-ADAPT-VQE is used. (b): $I_{4}$ as a function of number of SC-ADAPT-VQE layers used in the ansatz for a $L=8$ system. Each layer uses a longer-range operator $O_{d}$ given in Eq. 23 to create correlations that span more lattice sites. (c): An example of the exponential convergence of one of the angles parameterizing the $\lambda=0$ vacuum preparation circuit with one layer of SC-ADAPT-VQE, as a function of system size. The remainder of the angles are given numerically in App. F.
@@ IV.2 Wavepacket creation
The creation of particles on top of the prepared vacuum requires initializing local excited states. JLP proposed a method of creating wavepackets in the free theory by applying unitary evolution under $a_{k}^{\dagger}$ and $a_{k}$ defined in Eq. (3), and then adiabatically turning on the interaction as described in Sec. II. [^13] Whereas JLP specify the form of their wavepackets in position space, the method used in this work specifies the target profile in $k$-space. See App. B for a comparison. Instead of this, SVC is used to compress the wavepacket preparation circuit by taking advantage of classical computation. This step of the algorithm prepares a wavepacket whose size remains fixed as the total system size increases. While this method does not allow for the scalable preparation of wavepackets of increasing size (as will be necessary for simulations approaching the continuum), it utilizes the scalability in total system size, including vacuum regions. [^14] In this case, the classical computing capability limits the maximum size of the wavepacket. This is useful because a finer spacing in $k$-space increases the spatial momentum resolution of the wavepacket, building a more accurate representation of wavepackets in the continuum. Furthermore, wavepackets may only scatter between different lattice momenta in a lattice quantum field theory. Increasing the system size increases total number of lattice momenta and thus more faithfully reproduces the target continuum process.
@@
These are plotted as the dashed and faint solid lines respectively in Fig. 5b. $E_{k}$ and $v_{k}$ can be computed numerically in the digitized theory for any value of $\lambda$ by using ED and projecting onto each momentum sector. These numerically-calculated values are shown in Figs. 5a and b by the solid lines with circle and square markers for $L=10$.
@@
Figure 5: (a): The dispersion relation $E_{k}$ and (b): the group velocity $v_{k}$ for $m=1/2$ and $\lambda=0,2$. Exact computation is possible in the continuum (faint solid lines) and in the free lattice theory (dashed lines). The values of $E_{k}$ and $v_{k}$ are computed numerically by using ED on the Hamiltonian of Eq. (1) and projecting onto each momentum sectors (solid lines with circle and square markers). The numerical computations are done on a lattice of $L=10$ spatial sites with $n_{q}=2$ qubits per site. Discrete gradients are used for the numerical computations of $v_{k}$. (c): The convergence of the prepared particle wavepackets occupying three spatial sites with momentum $k=-\pi/3$, as a function of number of layers of the circuit from Fig. 3c. The local infidelity spanning the width of the wavepacket, $I_{3}$, is used to measure the quality of the prepared wavefunction for $\lambda=0,2$.
To simulate a high-energy collision, the highest momenta should be selected where the (digitized) lattice dispersion relation is still close to that of the continuum. With $L=10$ spatial sites, the resolution is poor and the digitized $v_{k}$ only match the continuum values at small $k$. The agreement can be improved with finer resolution by increasing the number of lattice sites. Finite-volume effects are reduced by using many lattice sites outside the interaction region. Based on this and on considerations of circuit depth and spread in $k$-space, wavepackets with $k=\pm\pi/3$ are chosen, with spread $\sigma_{k}=\pi/3$. The choice of $\sigma_{k}$ determines the support of the wavepacket in both position and momentum space. To limit the quantum resources required, the wavepacket width is truncated so that the circuits that create these states only act where the wavepackets have nonvanishing support. In this work, the spatial extent of the wavepackets is fixed to three lattice sites, and the circuits that prepare them act over $3n_{q}=6$ qubits. [^15] Similar to the field digitization, choosing the extent of the wavepackets in position and momentum space according to the Nyquist-Shannon Sampling Theorem will ensure a double-exponential convergence in precision for future calculations. A final choice in the setup of the simulation is the initial location of the wavepackets. The wavepackets should be well-separated at $t=0$ so they can be approximately treated as asymptotic “in” states. While this is strictly true only for infinite separation, the wavepackets are local, and their extent is fixed by the width of the circuit preparing them. Taking advantage of the locality of the $\phi^{4}$ interaction, the wavepackets are initialized to be separated by one spatial site. This construction may need to be modified for theories where interactions have less locality.
@@
Leveraging the analytic knowledge of the free theory, the wavefunction of the wavepacket $|\psi_{\text{wp}}\rangle$ defined in Eq. (10) is used as $|\psi_{\text{targ}}\rangle$ for the variational optimization. The adiabatic turn-on of interactions given by Eq. (12) is simulated classically, producing interacting wavepacket wavefunctions that are then used for $|\psi_{\text{targ}}\rangle$ in SVC for the interacting theory. See App. B for details on the classical determination of states used as $|\psi_{\text{targ}}\rangle$ for this step of state preparation. With this method, interacting wavepackets are prepared directly on top of the interacting vacuum. This way, in addition to reducing the circuit depth, SVC also removes the necessity for implementing the adiabatic evolution to turn on interactions.
@@
Because the region where the desired wavepacket has support does not grow with increasing system size $L$, the amount of qubits over which the variational circuit has to act is fixed by the spatial width of the wavepacket. Scaling these circuits is simple and straightforward since the number of parameters and structure do not change with system size. This is another example of using physical properties of the state, in this case locality of the wavepacket, to guide the design of the variational circuit. One layer of the brickwall ansatz used for this step of the state preparation is shown in Fig. 3c. The building blocks of this circuit are the variational gates defined in Ref. [140]. This block is particularly useful because any product of gates from the universal gate set $\{\text{CNOT},R_{X},R_{Y},R_{Z}\}$ can be written in terms of products of the block with different parameters. Each layer of the brickwall ansatz has a two-qubit gate depth of 2 and uses 20 variational parameters. [^16] The number of parameters can be reduced by incorporating the physical structure of the wavepackets (e.g., the spatial symmetry about the peak) into the circuits.
@@
Figure 5c shows the convergence of the prepared wavepacket state as a function of the number of layers of the brickwall ansatz of Fig. 3c. It is seen that the local infidelity spanning the wavepacket width, $I_{3}$, is systematically improved by adding more layers. Four layers of the circuit shown in Fig. 3c are used to prepare the free and interacting wavepackets. [^17] Note that with the parameters set to zero, inserting CNOTs between layers of this circuit will collapse the circuit to the identity. This is used in vacuum simulations to maintain the structure of the circuit. These brickwall circuits have two-qubit gate depths of 14 (including extra site-wise SWAP gates), and 80 parameters in total. While it is possible to run an extrapolation of the wavepacket parameters, this is not necessary. The angles parameterizing the vacuum creation circuits are observed to change on the scale of $10^{-4}-10^{-3}$ for L=1012L=10-12 (see Table 6), which is at the precision tolerance of IBM devices. The extrapolated angles for the vacuum preparation are used to prepare the equivalent vacuum state on a smaller system size for purposes of training the wavepacket circuit. The vacuum preparation angles extrapolated to $L=60$ are used in an $L=12$ system to optimize the wavepacket preparation parameters. The resulting parameters $\vec{\theta}_{\text{opt}}$ are given in Table 7, and numerical $I_{3}$ values are given in Table 1.
@@ IV.3 Time evolution
In this work, a translationally-invariant brickwall ansatz is used to compress the time evolution operator. Note that while there is no guarantee that a shallow brickwall circuit can approximate the time evolution operator well, time evolution under any Hamiltonian with local interactions is subject to Lieb-Robinson bounds [157]. Moreover, there is a maximum value for the group velocity of excitations from the lattice dispersion relation (Eq. (24)) [54] so it is reasonable to expect that low-depth brickwall circuits can reliably compress evolution under the Hamiltonian of Eq. (1) for sufficiently early times. The ansatz is made up of layers of the circuit shown in Fig. 3e, arranged in a symmetric way, seen in Fig. 6. Each variational step is made up of six forward layers and six reverse layers, with the reverse layers having the same parameters as the forward layers. This structure is chosen to mimic the structure of a second-order Trotter step. Each layer uses 12 parameters and has a two-qubit gate depth of two. The time evolution of the two-wavepacket state $e^{-itH}|\psi_{\text{2wp}}\rangle$ with 100 second-order Trotter steps on an $L=12$ system is used as $|\psi_{\text{targ}}\rangle$ to determine the variational circuits.
@@
Similar to the wavepacket preparation step, the $L$-extrapolation is not necessary to implement time evolution for the times simulated in this work. However, in this case this simplification is only valid for early times $(t\sim L)$, because the time evolution circuits act on the whole lattice. Since the vacuum is not prepared exactly, and the time evolution is not implemented exactly, the vacuum regions will evolve in time [20]. This evolution will be captured by the variational optimization procedure, up to finite-size effects that scale as O(t/L)O(t/L). [^18] These effects are estimated and the vacuum evolution is used for error mitigation. See App. E for details. To get around this limitation, the optimization can be run on larger systems for later times, or an $L$-extrapolation can be implemented. Propagating particles in a scattering simulation show more drastic effects from finite-size effects, such as wrapping around the lattice due to PBCs and re-scattering. Furthermore, larger wavepackets (needed for high-energy simulations or simulations approaching continuum) require more time to overlap, scatter, and fully separate. In these situations, it would be necessary to run the optimization for later simulation times on a larger lattice. While this is a limitation of the method in general, the scattering simulations are set up to require minimal time before scattering is observed (see Sec. IV.2 for details). This places a limit both on the time required in the simulation, as well as the size of the interaction region that needs to be used for training the variational circuits.
@@
One promising way to use variational circuits for time evolution without running into these finite-size effects is to compress a single time evolution step, and use that step repeatedly in place of a Trotter step. Similar methods have been investigated in Refs. [105, 110]. While the resulting circuits will be deeper than those produced by the current method, they will be more scalable with $L$. Another interesting direction is using MPS simulations of the time evolution operator for training [151, 110, 154]. This would remove the limitations placed on this method by finite-size effects. In this setup, the time evolution of one or many states may be used for training, as described in Sec. III. Methods that target time evolution operators in specific subspaces, such as restricted to the low-energy subspace, are currently being developed [152, 114]. The combination of subspace methods with variational time evolution has the potential to simplify the optimization task while also improving the quality of the produced circuits.
@@
Figure 7 shows the convergence of the time evolution variational circuits, compared to second-order Trotter circuits as a function of circuit depth. The plots show the infidelity as a function of two-qubit gate depth for the time evolution of a single wavepacket on an $L=6$ system. The Trotter time evolution circuits are constructed by adding more Trotter steps as described in Sec. II.1. Variational circuits fix the number of variational steps to two while increasing the depth of each step by adding layers of the circuit elements in Fig. 3d following the structure of Fig. 6. The results for t=19t=1-9 show that the brickwall circuits are able to consistently compress the Trotter circuits. This can especially be seen at late times, where considerably deeper Trotter circuits are required to accurately implement the time evolution operator, whereas the variational circuits show a modest growth in depth. Sharp dips in the curves of both the variational and Trotter infidelities are results of finite-size effects, as well as of numerical optimization (for the variational lines). A difference in the infidelity between $\lambda=0$ (top row) and $\lambda=2$ (bottom row) can be seen, which is attributed to using fixed-depth circuits to describe states with different correlation lengths. These plots show that by increasing the number of layers (and so, the number of parameters), the variational ansatz is able to represent the time-evolved state with increasing accuracy.
@@
Figure 7: The convergence of the time-evolved single-wavepacket state on an $L=6$ system with $\lambda=0$ (top row) and $\lambda=2$ (bottom row) for t=19t=1-9. The second-order Trotter time evolution (circles) is compared to the variational steps (squares) as a function of two-qubit gate depth. Two variational steps consisting of an increasing number of layers shown in Fig. 3d are used. The infidelity of the whole $L=6$ system with the exact time-evolved state is used to measure the quality of the wavepacket state as it propagates through the lattice.
@@
For the runs on the device, the number of variational steps is increased with the simulation time: $t=\{1,2,3\}$ use one variational step, $t=\{4,5,6\}$ use two, and $t=\{7,8,9\}$ use three. The number of layers in each step is kept fixed to six forward and six reverse layers, as in Fig. 6. The steps are optimized separately for each simulation time to find better quality solutions for the variational circuit. They are applied in a Trotter-like fashion, so that the number of parameters is the same for all time steps. The time evolution portion of the variational circuit uses 72 parameters in total. The parameters $\vec{\theta}_{\text{opt}}$ that implement the time evolution are given in Tables 8-16 and the local infidelities $I_{10}$ for each time step are shown in Table 1.